# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ 5.2.3. Adversarial Filtering
where $\theta_{\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}}$ are the kernels’ weights and $\delta$ is a modifier that directly optimizes $x_{w}$, similar to eq. (2). We maximize the spectral loss by sequentially propagating the input through a network of filters. Recall that direct modifications are not effective for disrupting low frequencies in the absence of filters due to their inclination to cause high-frequency disturbances only rather than systemic shifts in the values of many neighboring pixels in a uniform or gradual manner— see §5.2. Thanks to their convolutional nature, this no longer holds with our filters, making their outputs combinations of neighboring pixels. We still include $\delta$ that directly manipulates the input despite its inefficacy in the absence of filters. This is because the originally high-frequency disturbances caused by such modifications will now be distributed over neighbors of the receiving pixels due to our filters, translating them into useful low-frequency noise, causing further disruptions beyond the filters’ independent ability. Finally, we employ multiple ($T$) filters stacked after one another. The reason is that chaining several filters with various kernel sizes (geometries) targets different directions and slopes where the watermark could have been injected since watermarking is content-dependent and not all images can be changed similarly.
5.2.4. Spectral Loss
The former DFL is unsuitable for low frequencies (see §5.2) as it fails to force the changes to spread throughout the image. The collective energy at higher frequencies is much smaller in comparison, significantly reducing the total number of pixels that need to change to erase watermarks in these bands, which is achievable without violating the perceptual distance constraints. In contrast, sufficiently manipulating lower frequencies can rapidly change the structure, hindering convergence to an optimal solution if the distortion budget is not carefully distributed. This loss should also ensure our kernels indeed perform the filtering operation, which DFL fails to enforce. Better alternatives are:
Filter Regularization Loss (FRL): This loss enforces that our kernels perform the desired “smoothing” (filtering) and is given as:
$\begin{split}&\forall x\in\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}:\;FRL(x)=-\sum_{t\in[\![T]\!]}FRL_{\;t}(x)\\[-2.0pt] &\;\;\;\;\;\;\;\;\;\;\;s.t.\;\forall t:\;\;FRL_{\;t}(x)=\|x-\mu_{\frac{1}{2}}(x)^{{}^{\tiny\mbox{$M_{t}{\times}N_{t}$}}}\|_{2}\end{split}$
where $M_{t}\times N_{t}$ is the size of our $t^{th}$ kernel ${\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}}$ and $\mu_{\frac{1}{2}}(x)^{{}^{\tiny\mbox{$M_{t}{\times}N_{t}$}}}$ computes, for each pixel, the median value of this $M_{t}{\times}N_{t}$ vicinity around it. Maximizing FRL minimizes $\textit{\small\mbox{FRL}}_{\;t}$ for each filter $t$, which is the sum of all pixels’ distances from the median values in the filter’s environment around them in the output, bringing them as close as possible to these local medians. This eliminates variations to the extent possible under the perceptual loss, leading to smoother outputs. As the watermark should be distributed throughout the image, each patch will exhibit increased yet gradual “watermarking” variations that will be lost when bringing the values closer to these medians. A major benefit of FRL is also output consistency.
Mean Pool Loss (MPL): FRL is insufficient, as it only ensures proximity of the pixels to the local medians but imposes no constraints on these medians. The means of the various patches in the watermarked image are highly influenced by the watermark that causes many pixels to change their values. Due to the perceptual loss that strives to retain information, the procedure may converge to a sub-optimal solution, skewing these medians toward the original means and only slightly affecting pixel intensities, thereby retaining many of the shifts associated with the watermark. Thus, we require MPL:
$\forall x,y\in\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}:\;\;\;\;\;MPL^{{}^{\tiny\mbox{$L{\times}P$}}}(x,\;y)=\|\overline{x}^{{}^{\tiny\mbox{$L{\times}P$}}}-\;\overline{y}^{{}^{\tiny\mbox{$L{\times}P$}}}\|_{1}$
$\overline{x}^{{}^{\tiny\mbox{$L{\times}P$}}}$ denotes mean pooling with an $L\times P$ window. MPL takes the means over each $L\times P$ patch in both inputs and calculates the $\ell_{1}$ difference. As most pixels do not encode watermark information, maximizing the means’ differences (w.r.t. the watermarked image) while demanding proximity to the medians (using FRL) primarily affects pixels located farther away (those that encode the watermark) from all others, bringing them closer to the rest and eliminating watermarking variations. Additionally, proper low-frequency loss must spread the changes over multiple regions (see §5.2.4). MPL does so by dividing the image into patches and demanding that each receive a portion of the change, implying the number of patches must be large. Yet, each must contain enough pixels to manipulate low frequencies. We choose ${\small\mbox{{MPL}}}^{{}^{\tiny\mbox{$5{\times}5$}}}$ as a compromise, balancing these requirements. Calculating MPL for all filter sizes similar to FRL is less desirable as when the sizes shift away from this sweet spot, one of the conditions is violated. Note that, without FRL, MPL is insufficient as it can trivially lead to all pixels in each patch receiving a similar modification $\epsilon$, simply changing the patch’s brightness. This behavior satisfies the desired perceptual similarity but leaves the spectrum unaffected as the pixel variation rates remain identical.
5.2.5. Final Low-Frequency Destruction Procedure
[…]
$x_{nw}=\;\underset{\{\theta_{\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}}\},\delta}{argmin}\;\left[\begin{array}[]{cc}c_{{}_{\tiny\mbox{$\ell_{DVL}$}}}\cdot ReLU_{{}_{\tiny\mbox{$\ell_{DVL}$}}}(x_{w},\;x_{w}^{{}^{OF_{{}_{\tiny\mbox{$T$}}},\delta}})+\lambda_{{}_{\tiny\mbox{$\|\|_{2}$}}}\cdot\|x_{w}^{{}^{OF_{{}_{\tiny\mbox{$T$}}},\delta}}-x_{w}\|_{2}\\[5.0pt] -MPL^{{}^{\tiny\mbox{$5{\times}5$}}}(x_{w},\;x_{w}^{{}^{OF_{{}_{\tiny\mbox{$T$}}},\delta}})-\lambda_{{}_{\tiny\mbox{$FRL$}}}\cdot FRL(x_{w}^{{}^{OF_{{}_{\tiny\mbox{$T$}}},\delta}})\end{array}\right]$
We maximize the low-frequency loss $\ell_{F}={\small\mbox{{MPL}}}^{{}^{\tiny\mbox{$5{\times}5$}}}+\lambda_{{}_{\tiny\mbox{$FRL$}}}\cdot{\small\mbox{{FRL}}}$ by minimizing its negation similar to §5.1. $ReLU_{{}_{\tiny\mbox{$\ell_{DVL}$}}}$ ensures visual similarity, keeping the DVL distance from $x_{w}$ below $t_{{}_{\tiny\mbox{$\ell_{DVL}$}}}$. The resulting changes are gradual, obtained by applying our filters, whose weights are learned to ensure this behavior due to our spectral loss, yielding optimal $\{\widehat{\theta}_{\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}}\},\widehat{\delta}$ s.t. the final solution is $x_{nw}={\tiny\mbox{$x_{w}^{{}^{\widehat{OF}_{{}_{\tiny\mbox{$T$}}},\widehat{\delta}}}$}}$, where $\widehat{OF}_{{}_{\tiny\mbox{$T$}}}$ denotes the filters applied using $\{\widehat{\theta}_{\mathcal{K}^{*}_{{}_{{}_{\;\;t}}}}\}$. As each kernel’s values should be non-negative and sum to $1$, we apply Softmax to the learned weights to obtain effective values in this range. $\lambda_{{}_{\tiny\mbox{$FRL$}}}$ balances FRL and MPL and is calibrated s.t. both are simultaneously optimized, while $\lambda_{{}_{\tiny\mbox{$\|\|_{2}$}}}$ limits the impact of the $\|\|_{2}$ regularization. $c_{{}_{\tiny\mbox{$\ell_{DVL}$}}}$ balances the spectral and visual losses and is determined via a binary search, as in §5.1. We found a single filter network architecture universally effective. Hyperparameter details are in A.3.
Discussion: In the absence of our filters, the modifications at each step are less systemic, allowing different pixels to receive significant updates in opposing directions. Thus, the process will converge to a sub-optimal solution since massive changes that a subset of the pixels receive will substantially increase the visual loss, preventing further progress. While our low-frequency loss discourages this behavior and encourages uniform updates, it remains lacking as it is constantly countered by the visual loss that strives to retain similarity from a human’s perspective, limiting the number of pixels allowed to change their values, deeming such structural modifications less desirable compared to high-frequency changes that affect a discrete set of pixels thereby remaining potentially unnoticeable. Our filters, however, force the procedure to prioritize collective updates, drastically improving its performance. Readers are referred to A.5 for empirical results that establish these benefits.
This also explains why this attack, while effective against semantic watermarks, is unsuitable for removing non-semantic ones (see §6.2.2). The reason is that our low-frequency loss and filters encourage low-magnitude, yet collective, modifications and prevent significant updates to specific pixels (as opposed to §5.1). Yet, non-semantic watermarks are injected in high frequencies and thus rely on large variations between neighboring pixels. As this process can only slightly affect such peaks, the watermark will be preserved.
6. Experiments
6.1. Setup
Schemes. We experiment with seven SOTA schemes listed in Table 1 that cover the categories discussed in §2. They were selected since they recently appeared in top venues, and several of them remain unbroken by a query-free black-box attacker. Among the generator-specific schemes, the evaluated (original) implementations of Yu1 (Yu et al., 2022), Yu2 (Yu et al., 2021), and PTW (Lukas and Kerschbaum, 2023) watermark GANs (Goodfellow et al., 2014), while StableSignature (Fernandez et al., 2023) and TRW (Wen et al., 2023) operate on LDMs (Rombach et al., 2022) (see A.1 for details on synthesis algorithms). We are not aware of semantic schemes that appeared in top venues other than those we consider, which have also been the focus of previous works (Zhao et al., 2023c; Saberi et al., 2024; An et al., 2024). As most recent schemes are generator-specific, we include five such algorithms compared to two general-purpose systems.
Table 1. The SOTA schemes we study.
      General-Purpose Generator-Specific
Semantic     StegaStamp (Tancik et al., 2020) TRW (Wen et al., 2023)
Non-Semantic     HiDDeN (Zhu et al., 2018) Yu1 (Yu et al., 2022), Yu2 (Yu et al., 2021), PTW (Lukas and Kerschbaum, 2023), StableSignature (Fernandez et al., 2023)
Data. Most generator-specific schemes require no data, generating random images. Yet, as TRW and StableSignature watermark text-to-image LDMs (StableDiffusion (hf, [n. d.])), we use prompts from Stable Diffusion’s test set. For HiDDeN, we use test samples from the COCO dataset (Lin et al., 2014) on which it was trained. StegaStamp watermarks $400{\times}400$ images. Thus, we downscale samples from the known CelebA-HQ-1024 (Karras et al., 2017) [^2] FFHQ-1024 (Karras et al., 2019) samples yield similar results. For brevity, we focus on CelebA-HQ.. We evaluate each system on $100$ random images, similar to previous work (Jiang et al., 2023).
Metrics. Removing a watermark entails satisfying two conditions: 1) It must no longer be detected, and 2) the output should retain quality and proximity to the watermarked image (see §3).
Detection: For binary watermarks, detection occurs if the ratio of bits matching the watermark in the sequence extracted from the image is above some threshold $t_{w}$. As in previous work (Zhao et al., 2023c; Lukas et al., 2024), we select $t_{w}$ that can reject the null hypothesis $H_{0}$ that $k$ or more matching bits were randomly extracted with $p<0.01$. This has probability $P(X\geq k|H_{0})=\sum_{i=k}^{n}{n\choose k}0.5^{n}$, for watermark length $n$. Thus, $t_{w}=\frac{k}{n}$ for which this probability is below $0.01$ is selected. We denote the portion of successfully extracted bits as the bit accuracy, which we compare against $t_{w}$. Watermark lengths are in A.2. TRW’s watermarks are unbounded, and detection occurs if the $\ell_{1}$ distance— or effectively, the Mean Absolute Error (MAE)— of the extracted sequence from the watermark is below a threshold determined s.t. the ratio of falsely detected non-watermarked images ($FPR$) is low. We chose a higher $FPR$ ($0.02$) than the original paper (Wen et al., 2023) ($0.01$), making the attacker’s task harder, arriving at $MAE=71$. For consistency with other schemes, we use this distance’s inverse: for the watermark to be detected, the $MAE$’s inverse must be $\geq t_{w}=\frac{1}{71}$.
Quality: We use two metrics common in previous work (Zhao et al., 2023c; An et al., 2024; Lukas et al., 2024; Hu et al., 2024; Lukas and Kerschbaum, 2023; Fernandez et al., 2023): Learned Perceptual Similarity (LPIPS (Zhang et al., 2018)) and Frèchet Inception Distance (FID)  (Heusel et al., 2017). To eliminate biases, we refrain from using the LPIPS methods UnMarker employs (see §5.1) as our attacks explicitly optimize these metrics. Instead, we use well-known LPIPS-VGG (Zhang et al., 2018). FID evaluates the attack samples’ distributions w.r.t. the datasets of watermarked images, measuring dataset-wise similarity.
Baselines. We include various typical manipulations. Their parameters were chosen following previous work (Lukas and Kerschbaum, 2023; Zhao et al., 2023c) and to allow maximum distortion while retaining acceptable quality. These are Mild Center Cropping, JPEG compression, Quantization, and Gaussian Blurring (filtering) with a $5{\times}5$ kernel. We also evaluate Guided Blurring to demonstrate how even these stronger filters cannot break robust watermarks, as argued in §5.2. Their kernel size is also $5{\times}5$, but we exploit their edge-preserving properties to increase its std. (to $5.0$). To control the modulation at edges, these accept an additional color kernel’s std. (see §5.2), for which we have three values: 0.1, 1.00.11.00.1,\;1.00.1 , 1.0, and $5.0$. Yet, $1.0$ and $5.0$ are overly destructive and do not preserve quality, chosen only to better explore robustness. We also evaluate UnMarker against the SOTA DiffusionAttack (Zhao et al., 2023c; Saberi et al., 2024; An et al., 2024) and VAEAttack (Zhao et al., 2023c; An et al., 2024). To our knowledge, these regeneration techniques are the only existing practical (black-box and query-free) and effective attacks, as adversarial attacks make unrealistic relaxing assumptions (see §2) and are therefore disqualified.
6.2. Evaluations
6.2.1. Watermark Utility
First, we ensure the chosen schemes’ utility, following the requirements from §4.1: Applicability, Stealthiness, and Robustness to Manipulations. While the latter demands withstanding basic manipulations, true robustness is contingent on resisting advanced attacks as well. We separate these cases for consistency with watermark papers that mainly consider standard manipulations. Stronger attacks are in §6.2.3. We focus on Applicability and Robustness since we use original implementations of SOTA systems whose stealthiness has been shown by their authors.
Visual metrics are excluded here, as evaluating the naive manipulations’ output quality is not our objective, knowing that our basic settings are quality-preserving (Zhao et al., 2023c; Lukas and Kerschbaum, 2023). We do not evaluate the rotation manipulations by Zhao et al. (Zhao et al., 2023c), following their arguments w.r.t. their impracticality as they can be easily spotted and corrected before invoking the detector. Still, their finding of some of the more robust schemes being vulnerable to rotations may seem to contradict our assertions in §4.1 that these watermarks are in the spectral amplitudes as such watermarks should remain unaffected by rotations that preserve these magnitudes. Indeed, some schemes were found agnostic to rotations. As for the systems exhibiting this vulnerability, note that watermarking schemes are generally not explicitly designed to construct watermarks in the spectral amplitudes. Rather, this behavior is naturally adopted when trained for robustness to various practical manipulations (see §4). Failure to expose these systems to a wider set of demands during training (e.g., resistance to rotations) may result in them mistakenly associating additional indicators, such as a specific image orientation, with the watermark. Yet, as rotations are impractical, this is not an issue.
Table 2. Performance of watermarking schemes against each stage of UnMarker. The complete attack is in the last row (CHL). Numbers in parentheses are the detection thresholds, while Detect is the portion of images meeting them.
      Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSignature (69%) StegaStamp (63%) TRW (0.0141)
[…]
@@ 6.2.1. Watermark Utility
A detailed analysis is deferred to A.4 alongside the results (Table 5). All schemes behave as desired without manipulations, detecting watermarked images: For binary watermarks, the bit accuracies are far above the thresholds, approaching $100\%$. For TRW, the inverse distance is $0.022$, well over the $0.0141$ threshold. All systems are fairly resistant to manipulations. Semantic schemes remain almost unaffected. Among non-semantic schemes, Yu1 and StableSignature are the most resistant, retaining robustness against all realistic configurations. Yet, the latter shows some vulnerability to Guided Blurring with overly destructive parameters ($color\;std.\geq 1.0$), but since this is not a practical attack, it remains reliable. PTW rejects most manipulations but is affected by reasonable Quantization ($10$), while the less advanced HiDDeN is vulnerable to JPEG and Blurring. Yu2 resists all manipulations except cropping, indicating high dependency on spatial attributes, making it unsuitable for watermarking, as it violates the robustness requirement from §4.1.
@@
The results demonstrate how all schemes aside from Yu2 are robust to spatial manipulations such as cropping, indicating their watermarks reside in a different channel, which we know and will next verify empirically is the spectral amplitudes. This is evident for PTW and HiDDeN that exhibit vulnerability to Quantization and JPEG that affect high frequencies. The resistance of other schemes to these manipulations, however, does not imply robustness but rather that their watermarks are harder to remove. Yet, advanced attackers can still break these systems, as we show below. Yu2 is highly dependent on information encoded in the background’s pixels, suggesting its watermarks are not constructed as robust, collective measurements but rather in volatile spatial locations.
6.2.2. UnMarker

### v2

@@ 5.2.5 Final Low-Frequency Destruction Procedure
$x_{nw}=\;\underset{\{\theta_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}\},\delta}{argmin}\;\left[\begin{array}[]{cc}c_{{{\text{\tiny\mbox{${DVL}$}}}}}\cdot ReLU_{{{\text{\tiny\mbox{${DVL}$}}}}}(x_{w},\;x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}})+\lambda_{{\|\|_{2}}}\cdot\|x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}}-x_{w}\|_{2}\\[5.0pt] -MPL^{{}^{\text{\tiny\mbox{$5{\times}5$}}}}(x_{w},\;x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}})-\lambda_{{\text{\tiny\mbox{$FRL$}}}}\cdot FRL(x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}})\end{array}\right]$
We maximize the loss $\ell_{F}\!=\!\!{\text{{MPL}}}^{{}^{\text{\tiny\mbox{$5{\times}5$}}}}\!\!\!\!+\lambda_{{\text{\tiny\mbox{$FRL$}}}}\cdot{\text{{FRL}}}$ by minimizing its negation similar to §5.1. $ReLU_{{{\text{\tiny\mbox{${DVL}$}}}}}$ ensures similarity, keeping the DVL distance from $x_{w}$ below $t_{{{\text{\tiny\mbox{${DVL}$}}}}}$. The resulting changes are gradual due to our filters, whose weights are learned to achieve this behavior thanks to our spectral loss, yielding optimal $\{\widehat{\theta}_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}\},\widehat{\delta}$ s.t. the final solution is $x_{nw}\!\!=\!\!\text{\mbox{$x_{w}^{{}^{\widehat{OF}_{{}_{\text{\tiny\mbox{$T$}}}},\widehat{\delta}}}$}}\!\!\!\!$, where $\widehat{OF}_{{}_{\text{\tiny\mbox{$T$}}}}$ denotes the filters applied using $\{\widehat{\theta}_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}\}$. As each kernel’s values should be non-negative and sum to $1$, we apply Softmax to the learned weights to obtain effective values in this range. $\lambda_{{\text{\tiny\mbox{$FRL$}}}}$ balances FRL and MPL and is calibrated s.t. both are optimized, while $\lambda_{{\|\|_{2}}}$ limits the impact of the $\|\|_{2}$ regularization. $c_{{\text{\tiny\mbox{$DVL$}}}}$ balances the spectral and visual losses and is determined via a binary search, as in §5.1. We found a single filter network architecture universally effective. Details are in Appendix A.
Discussion: Without our filters, the modifications at each step are less systemic, allowing different pixels to receive significant updates in opposing directions. Thus, we will converge to a sub-optimal solution since massive changes that a subset of pixels receive will substantially increase the visual loss, preventing further progress. While our low-frequency loss discourages this behavior and encourages uniform updates, it remains lacking, constantly countered by the visual loss that strives to retain similarity, limiting the number of pixels allowed to change their values, deeming such structural modifications less desirable compared to high-frequency changes that affect discrete pixels thereby remaining potentially unnoticeable. Our filters force prioritizing collective updates, drastically improving performance. Results establishing these benefits are in Appendix D.
While effective against semantic watermarks, this attack is unsuitable for non-semantic ones (see §6.2.2); The low-frequency loss and filters encourage low-magnitude, yet collective, changes and prevent significant updates to specific pixels (as opposed to §5.1). Yet, non-semantic watermarks are injected in high frequencies and thus rely on large variations between neighboring pixels. As this process can only slightly affect such peaks, the watermark is preserved.
6 Experiments
6.1 Setup
Schemes. We consider the SOTA watermarking schemes in Table I, covering all categories from §2, selected for appearing in top venues, with several remaining unbroken by a query-free black-box attacker. Among the generator-specific schemes, Yu1 [21], Yu2 [22], and PTW [23] watermark Generative Adversarial Networks (GANS) [53], while StableSignature [20] and TRW [18] operate on LDMs [46]. We are not aware of semantic schemes from top venues other than those we consider, which are the focus of previous work [35, 29, 31]. As most recent systems are generator-specific, we include five such schemes and two general-purpose ones. Compared to existing works (see §2.2), we span a wider variety of recent systems.
TABLE I: Studied SOTA watermarking schemes.
      General-Purpose Generator-Specific
Semantic     StegaStamp [24] TRW [18]
Non-Semantic     HiDDeN [19] Yu1 [21], Yu2 [22], PTW [23], StableSignature [20]
Data. Most generator-specific schemes require no input data. They were trained on various datasets: TRW and StableSignature on LAION-5B [54], Yu1 and Yu2 on CelebA [55], and PTW on FFHQ [56]. As TRW and StableSignature watermark text-to-image LDMs (StableDiffusion [2]), we use StableDiffusion’s prompts. For HiDDeN, we take samples from COCO [57] on which it was trained. StegaStamp watermarks $400{\times}400$ images. Thus, we downscale CelebA-HQ-1024 samples [^2] FFHQ-1024 samples yield similar results (excluded for brevity).. We evaluate each system on $100$ random images, similar to previous work [30].
Metrics. Following §3, we are concerned with two metrics:
1) Detection: Bit Accuracy. For binary watermarks, detection occurs if the ratio of bits matching the watermark in the sequence extracted from the image is above some threshold $t_{w}$. As in previous work [35, 32], we select $t_{w}$ that can reject the null hypothesis $H_{0}$ that $k$ or more matching bits were randomly extracted with $p<0.01$. This has probability $P(X\geq k|H_{0})=\sum_{i=k}^{n}\binom{n}{k}\frac{1}{2^{n}}$, for watermark length $n$. Thus, $t_{w}=\frac{k}{n}$ for which this probability is below $0.01$ is selected. We denote the portion of successfully extracted bits as the bit accuracy, which we compare against $t_{w}$. Watermark lengths are in Appendix B.
Inverse Distance. TRW’s watermarks are unbounded, and detection occurs if the $\ell_{1}$ distance (effectively, the Mean Absolute Error— MAE) of the extracted sequence from the watermark is below $t_{w}$ determined s.t. the ratio of falsely detected non-watermarked images (FPR) is low. We chose a higher FPR ($0.02$) than the original paper [18] ($0.01$), making the attacker’s task harder, obtaining $\textit{\text{MAE}}=71$. For consistency, we use this distance’s inverse: for the watermark to be detected, the MAE’s inverse must be $\geq t_{w}=\frac{1}{71}$.
Security Advantage. This binary indicator is the scheme’s ability to detect over $50\%$ of watermarked images. Most genuine content is not controversial, and detectors will not be invoked to determine its authenticity. They are only necessary when suspicious content surfaces. Absent additional knowledge, we assume a natural-looking controversial image has equal probabilities of being real or fake. Here, the best trivial detector that achieves the highest average accuracy on suspicious images would be a random classifier. An attacker will have a $50\%$ chance of fooling it with a fake image. For a scheme to have a Security Advantage, it must detect watermarked images at a higher rate. Otherwise, the resources invested in it will be in vain as it offers no additional protection against attackers compared to random systems. While such a scheme will perform better on non-watermarked images as it is trained to detect the existence of a watermark, this is irrelevant from a security aspect where only identifying the fake (watermarked) content matters.
2) Quality: We use two common  [35, 31, 32, 33, 23] human vision-inspired metrics: LPIPS [51] and Frèchet Inception Distance (FID) [58]. For fairness, we avoid LPIPS methods UnMarker explicitly optimizes (see §5.1), resorting to LPIPS-VGG [51]. FID compares the attack samples’ distribution to that of watermarked images, measuring dataset-wise similarity. We exclude geometric metrics (PSNR or norms) as we find them misleading. For instance, although visual metrics (Table III) and human inspection (Appendix E) verify the VAEAttack’s inferior quality, geometric methods deem it comparable with UnMarker.
Basic Manipulations. We evaluate Center Cropping ($10\%$), JPEG, Quantization, and Gaussian Blurring (filtering). Their parameters (see Appendix C) were chosen following previous work [23, 35] and to allow maximum distortion while retaining acceptable quality. We also evaluate Guided Blurring, showing how even these stronger filters cannot break robust watermarks, as argued in §5.2. To better explore these advanced filters, we even evaluate them with overly destructive parameters that do not represent viable attacks.
SOTA Attacks. We compare UnMarker against SOTA watermark removal attacks: The DiffusionAttack [35, 29, 31] and VAEAttack [35, 31]. To our knowledge, these regeneration methods are the only practical (query-free and black-box) and effective attacks, as adversarial attacks are disqualified for their unrealistic assumptions (see §2).
6.2 Evaluations
6.2.1 Watermark Utility
We ensure the chosen schemes’ utility, following the requirements from §4.1: Applicability, Stealthiness, and Robustness to Manipulations. While the latter demands withstanding basic manipulations, true robustness is contingent on resisting advanced attacks as well. We separate these cases for consistency with watermark papers that mainly consider basic manipulations. Advanced attacks are in §6.2.3. We focus on Applicability and Robustness as we use original implementations of SOTA systems whose Stealthiness has been shown by their authors.
Visual metrics are excluded here, as evaluating the naive manipulations’ output quality is not our objective, knowing that our basic settings are quality-preserving [35, 23]. We do not consider rotation manipulations, following Zhao et al.’s [35] arguments w.r.t. their impracticality as they can be easily spotted and corrected before invoking the detector. Still, their finding of some robust schemes vulnerable to rotations may seem at odds with our assertions that these watermarks are in the collective spectral amplitudes, which should leave them unaffected by rotations. Yet, schemes are generally not explicitly designed to construct watermarks in the spectral amplitudes. Rather, this behavior is adopted when trained for robustness (see §4). Failure to expose these systems to a wider set of demands (e.g., resistance to rotations) may result in them mistakenly associating additional indicators, such as a specific orientation, with the watermark. As rotations are impractical, this is not an issue.
TABLE II: Performance of watermarking schemes against each stage of UnMarker.
Note: Performance without attacks and under cropping is included for reference. The complete attack is in the last row (CHL). Numbers in parentheses are
the detection thresholds, while Detect is the portion of images meeting them. The best attacks are in bold font. All schemes lose their Security Advantage.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSig. (69%) StegaStamp (63%) TRW (0.0141)
@@ 6.2.1 Watermark Utility
Cropping (10%)    74.56% 100% 53.71% 0% 86.84% 100% 93.9% 100% 97.86% 100% 93.96% 100% 0.0155 100%
@@
A detailed analysis is in Appendix C alongside the results (Table IV). All schemes detect watermarked images without manipulations: For binary watermarks, all bit accuracies approach $100\%$. For TRW, the inverse distance is $0.022$, well over the $0.0141$ threshold. All systems are fairly resistant to manipulations. Semantic schemes remain almost unaffected. Among non-semantic schemes, Yu1 and StableSignature are the most resistant, retaining robustness against all realistic configurations. The latter shows some vulnerability to Guided Blurring with overly destructive parameters, but since this is not a practical attack, it remains reliable. PTW rejects most manipulations but is affected by reasonable Quantization ($10$), while the less advanced HiDDeN is vulnerable to JPEG and Blurring. Yu2 resists manipulations except cropping, indicating high dependency on spatial attributes, making it unsuitable for watermarking, as it violates the robustness requirement from §4.1.
@@
The results demonstrate how all schemes except Yu2 are robust to spatial manipulations (e.g., cropping), indicating their watermarks are in a different channel, which we next verify is the spectral amplitudes as expected. This is evident for PTW and HiDDeN that exhibit vulnerability to Quantization and JPEG that affect high frequencies. The resistance of other schemes to these manipulations merely implies that their watermarks are harder to remove. Yet, advanced attacks can still break these systems, as shown below. Yu2 highly depends on information in the background, suggesting its watermarks are not constructed as robust, collective measurements but rather in volatile spatial locations.
6.2.2 UnMarker